3.Using the definition of "Big-O" determine if each of the following functions, f(x)=(xlogx)^2−4 and g(x)=5x^5 are O(x^4) and prove your claims.


1
Expert's answer
2021-11-08T21:20:07-0500

Definition: Let "f" and "g" be real-valued functions. We say that "f(x)" is "O(g(x))" is there exists two numbers "C" and "k" such that "|f(x)| ≤ C|g(x)| \forall x > k." or in limit terms "\lim\limits_{x\to+\infty}|\frac{f(x)}{g(x)}|<C."

Let's consider "\lim\limits_{x\to+\infty}\frac{(xlogx)^2−4}{x^4}=\lim\limits_{x\to+\infty}\frac{log^2x}{x^2}-\lim\limits_{x\to+\infty}\frac{4}{x^4}=0" . So "f(x)=(xlogx)^2\u22124" is "o(x^4)."

Let's consider "\lim\limits_{x\to+\infty}\frac{5x^5}{x^4}=\lim\limits_{x\to+\infty}5x=+\infty." So for all C "\exist k |5x^5|>C |x^4|" for all "x>k" . So, "g(x)=5x^5" is not an "O(x^4)" .